# The Security Threat of Compressed Projectors in Large Vision-Language Models

### v1

@@ Abstract
The choice of a suitable visual language projector (VLP) is critical to the successful training of large visual language models (LVLMs). Mainstream VLPs can be broadly categorized into compressed and uncompressed projectors, and each offering distinct advantages in performance and computational efficiency. However, their security implications have not been thoroughly examined. Our comprehensive evaluation reveals significant differences in their security profiles: compressed projectors exhibit substantial vulnerabilities, allowing adversaries to successfully compromise LVLMs even with minimal knowledge of structural information. In stark contrast, uncompressed projectors demonstrate robust security properties and do not introduce additional vulnerabilities. These findings provide critical guidance for researchers in selecting optimal VLPs that enhance the security and reliability of visual language models. The code will be released.
The Security Threat of Compressed Projectors in Large Vision-Language Models
Yudong Zhang1,2, Ruobing Xie2, Jiansheng Chen3, Xingwu Sun2,4, Zhanhui Kang2, Di Wang2, Yu Wang1 1Tsinghua University, 2Large Language Model Department, Tencent, 3University of Science and Technology Beijing, 4University of Macau, zhangyd16@mails.tsinghua.edu.cn, xrbsnowing@163.com, jschen@ustb.edu.cn, sunxingwu01@gmail.com, kegokang@tencent.com, diwang@tencent.com, yu-wang@mail.tsinghua.edu.cn.
@@ 1 Introduction
Large vision-language models (LVLMs) have achieved remarkable success in various multimedia applications Dai et al. (2023); Liu et al. (2023a); Bai et al. (2025); Wang et al. (2024); Chen et al. (2024), particularly in tasks like visual question answering (VQA). A typical LVLM framework consists of three key components: visual encoder (VE), large language model (LLM), and vision-language projector (VLP). Current training methodologies primarily focus on optimizing the VLP to effectively integrate visual features extracted from a pre-trained VE into a pre-trained LLM. This approach substantially minimizes computational demands when compared to building an LVLM from the ground up, thereby allowing researchers to leverage existing advancements. Moreover, it ensures greater stability and reliability in both the training process and practical deployment of LVLMs.
Recently, a growing body of research has employed CLIP Radford et al. (2021); Fang et al. (2023) as the VE and initiated training starting from various pre-trained LLMs. The choice of VLP plays a crucial role in determining the effectiveness of LVLM training. VLPs can generally be classified into two distinct categories: compressed projectors and uncompressed projectors. Compressed projectors, exemplified by Q-former Li et al. (2023); Dai et al. (2023); Zhu et al. (2024), achieve high computational efficiency by compressing a larger number of visual tokens into a smaller dimension using query tokens. On the other hand, uncompressed projectors, typified by MLP-based architectures Liu et al. (2023a); Shi et al. (2024), convert visual tokens to feature dimensions matching the LLM, with the number of output tokens proportional to the number of input tokens.
Previous studies have tended to focus only on the security of visual encoders or LLMs. On the visual encoder, Wang et al. (2023); Yin et al. (2023) present adversaries with an opportunity to compromise model security by targeting solely the visual encoder. As for LLM, previous studies Carlini et al. (2023); Zou et al. (2023); Chao et al. (2023) have fully explored the safety of LLM, and the user has the option of choosing a better performing and safer LLM on his or her own. However, the security of visual language projectors has not been fully explored, only few studies have focused on attacks on VLPs. Although some studies Zhang et al. (2025) have attempted to attack the VLP (mainly Q-former), they have not thoroughly analyzed the security of the VLP structure.
@@
We assessing the vulnerabilities of both compressed and uncompressed projectors by conducting adversarial attacks across diverse white-box and gray-box scenarios to evaluate the security implications for different VLP architectures. Our findings reveal three key insights: (1) Compressed projectors exacerbate security vulnerabilities beyond those affecting visual encoders. Specifically, an adversary can significantly degrade model performance by targeting compressed projectors, even with limited knowledge about the VLP. (2) In contrast, uncompressed projectors do not introduce additional security risks, as attacking them yields results comparable to attacking visual encoders directly. (3) This discrepancy in security between compressed and uncompressed projectors stems from their architectural differences and remains independent of the number of visual tokens. Notably, even when the visual tokens of an uncompressed projector are reduced (e.g., through pooling) to approach those of a compressed projector, the uncompressed projector still maintains its robust security characteristics.
Based on our analyses, we propose the following suggestions: (1) We strongly recommend employing uncompressed projectors rather than compressed projectors in high-security environments to mitigate potential risks from adversarial attacks. (2) In scenarios where computational efficiency is a priority, researchers can implement techniques like pooling operations to proportionally reduce the number of output tokens generated by uncompressed projectors. This approach offers enhanced security compared to using compressed projectors while maintaining acceptable performance levels, thereby striking an appropriate balance among effectiveness, efficiency, and security.
Our research provides three key contributions: (1) This study is the first to systematically investigate VLP security, presenting a comprehensive and systematic evaluation of various VLP architectures. (2) Our experimental results uncover significant security vulnerabilities in compressed projectors, revealing critical implications for models that employ such components. (3) We demonstrate that operating on visual tokens generated by uncompressed projectors offers enhanced security compared to compressed projectors when the number of visual tokens are largely comparable, thereby identifying promising approaches for VLP selection under efficiency constraints.
@@ 2 Related Work
Large Vision-Language Models. Training LVLMs from scratch is often prohibitively expensive. Thus, current popular frameworks typically begin with a pre-trained unimodal visual encoder and a large language model, focusing their training efforts on developing a VLP to connect the two modalities and effectively incorporate visual features into the LLM framework. The mainstream VLPs currently fall into two primary categories: compressed projectors and uncompressed projectors. The compressed projectors, such as Q-former Li et al. (2023), Resampler Alayrac et al. (2022), and D-Abstractor Cha et al. (2024), reduce the number of visual tokens to a fixed number of output tokens. In contrast, uncompressed projectors, exemplified by MLP Liu et al. (2023a), maintain a proportional relationship between the number of visual tokens and their corresponding output tokens.
Compressed and uncompressed projectors present distinct trade-offs in terms of computational efficiency, memory usage, and implementation complexity. On one hand, uncompressed projectors offer simplicity in design but incur significant computational costs. Specifically, their computational requirements increase significantly as their length scales linearly with the square of the resolution Chen et al. (2024). Furthermore, in video processing applications Ren et al. (2024), this scaling behavior results in a linear growth of required length relative to the number of video frames. On the other hand, compressed projectors achieve improved efficiency by reducing redundancy within the visual space. This is accomplished through the compression of token lengths from the visual encoder to a specified capacity, enabling strong performance while maintaining high computational efficiency. While prior research has extensively studied the performance and efficiency trade-offs between these two approaches in VLPs, there remains a critical gap: the safety implications of these methods have been largely neglected.
@@
This paper focuses on a practical gray-box scenario where the adversary is aware of both the VE weights and the structural of VLP, whether it be compressed or uncompressed. By crafting specialized loss functions tailored to the VLP architecture, the adversary can effectively execute targeted adversarial attacks against LVLMs within both the white-box and gray-box framework. Our findings not only expose the vulnerabilities inherent in compressed projectors but also provide new insights into selecting between compressed and uncompressed projectors from a security perspective.
@@ 3.1 Preliminaries
Notations. We provide a brief overview of the definition of notations. A LVLM, denoted as $f$, generally comprises three key components: (1) the visual encoder $f_{\text{VE}}$, typically based on CLIP; (2) the vision-language projector $f_{\text{VLP}}$, which is usually implemented as a Q-former or an MLP; and (3) the large language model $f_{\text{LLM}}$. The LVLM $f$ accepts an image $x_{i}$ and an instruction $x_{t}$ as input, producing an output $y$. The process begins with the input image $x_{i}$ being processed by the $f_{\text{VE}}$ to generate the visual feature representation $f_{\text{VE}}(x_{i})$. Next, depending on whether it is a compressed or uncompressed VLP: (1) for the compressed VLP, $f_{\text{VLP}}$ extracts relevant features from both the visual features $f_{\text{VE}}(x_{i})$ and the instruction $x_{t}$; while (2) for the uncompressed VLP, $f_{\text{VLP}}$ transforms the visual token dimension to align with the LLM input space. Both processes result in the projection output $f_{\text{VLP}}(f_{\text{VE}}(x_{i}),x_{t})$. These projected features are then fed into $f_{\text{LLM}}$, which generates the final output $y$ as Eq. 1:
@@
Compressed projectors. The Q-former stands as the canonical representative of compression projectors, functioning as a lightweight yet trainable querying transformer. It effectively extracts textually relevant features from the VE CLIP through a set of learnable queries. Notably, during BLIP-2’s two-stage pre-training process, both the VE and the LLM remain static, with only the Q-former and its associated queries being trainable components. This training paradigm has been successfully employed in subsequent studies, such as InstructBLIP Dai et al. (2023) and MiniGPT-4 Zhu et al. (2024), which also adopt similar approaches.
Uncompressed projectors. A typical approach involves using a simple multilayer perceptron (MLP) to project the visual token representation into the input space of a LLM. For instance, in LLaVA v1.5 Liu et al. (2023a), the visual encoder outputs a 1024-dimensional embedding, which is then mapped to 4096 dimensions via an MLP to align with the input requirements of the Vicuna LLM. Similarly, Eagle Shi et al. (2024) concatenates feature representations from multiple visual encoders to form a longer vector (typically ranging from 6000 to 8000 dimensions, depending on the number of visual encoders used), which is subsequently projected down to 4096 dimensions using an MLP.
@@ 3.2 An Empirical Analysis of Component Accessibility Within LVLM Attacks
Vision-language Projector (VLP). We assume the adversary possesses knowledge of the VLP’s architectural details, such as its implementation as a Q-former or MLP. Notably, all LVLMs built on Q-former, including BLIP-2, InstructBLIP, and MiniGPT-4, share identical VLP architectural designs. However, while the adversary may be aware of these structural specifics, obtaining the exact model parameters remains challenging due to the diversity in pre-trained LLMs and training datasets.
@@
The white-box setting. In the white-box setting, where adversaries have full access to model parameters (including both VEs and VLPs), we evaluated the safety of VLPs under worst-case scenarios. The adversary aims to modify a clean image to create an adversarial example capable of fooling the LVLM, leading to incorrect or unintended responses.
The gray-box setting. Building on the analysis presented, we focus on a practical yet hitherto unexplored gray-box setting, in which the adversary has access to the weights of the VE CLIP and the architectural structure of the VLP but lacks access to either the weights of the VLP or any information about the LLM. In this context, we can obtain a surrogate VLP model through transfer learning from other LVLMs, or alternatively, train the surrogate VLP model from scratch.
@@ 3.3 Surrogate Models Used in Attacks
Based on our analysis of the accessibility levels of various LVLM components presented in Sec. 3.2, we design and implement adversarial attacks against VLPs under three distinct scenarios with increasing levels of difficulty to comprehensively assess their security vulnerabilities. (1) The first scenario represents the simplest case, where the attack operates in a white-box setting, allowing full visibility into the model’s architecture and parameters. (2) The second scenario involves attacking a surrogate VLP from a similar LVLM, targeting another specific LVLM. (3) The most challenging scenario requires an attacker to develop an surrogate VLP from scratch, significantly increasing the complexity of generating effective adversarial examples.
@@
[…]
Surrogate models trained from scratch in gray-box setting. When similar LVLMs are unavailable for transfer, the structural insights into VLP provide us with a unique opportunity. Once the target model’s VLP structure (using compressed or uncompressed projectors) is identified, we can train surrogate Q-formers or MLPs from scratch. The detailed training procedure is outlined in Sec. 4.1. Notably, during this process, no information about the specific LLM used by the target LVLM is required; only the VE weights and VLP structure are necessary for training the surrogate VLP.
Figure 1: The attack pipeline operates by first extracting surrogate VLPs $\{f_{\text{VLP}}^{1},\dots,f_{\text{VLP}}^{N}\}$, which are then utilized to generate adversarial examples $x^{\prime}_{i}$ through the loss functions $\mathcal{L}_{\text{VE}}$, $\mathcal{L}_{\text{VLP}}$ and $\mathcal{L}_{\text{TCP}}(\beta,K)$. A key aspect of this investigation is determining whether incorporating attacks on VLPs increases security vulnerabilities compared to solely attacking VEs, thereby assessing the robustness of the VLP structure.
@@ 3.4 Loss Function for More Effective Attacks
To evaluate the security of the VLP structure, we employ adversarial attack methods to analyze the robustness of LVLM. For a model with loss function $\mathcal{L}$, Fast Gradient Sign Method (FGSM) Goodfellow et al. (2015) performs an adversarial attack through gradient ascent as follows: $x^{\prime}_{i}=x^{\prime}_{i}+\nabla_{x^{\prime}_{i}}\mathcal{L}$. In contrast, Projected Gradient Descent (PGD) Madry et al. (2018) conducts multiple iterations of gradient ascent and projects $x^{\prime}_{i}$ onto the constrained perturbation space after each step. Typically, PGD restricts the $l_{\infty}$ norm distance with $\|x_{i}-x^{\prime}_{i}\|_{\infty}\leq\epsilon$. Additionally, a variant of the Carlini & Wagner (C&W) attack Carlini and Wagner (2017), referred to as CW-$l_{2}$, aims to maximize the loss function $\mathcal{L}$ while minimizing the $l_{2}$ norm distance $||x_{i}-x^{\prime}_{i}||_{2}^{2}$.
To compare the security implications introduced by the VLP structure, we evaluate the attack results on VE and VLP. For a clean image input, let $V_{\text{out}}=f_{\text{VE}}(x_{i})$ and $P_{\text{out}}=P(f_{\text{VE}}(x_{i}),x_{t})$ denote the outputs of VE and VLP, respectively. For an adversarial image, let $V^{\prime}_{\text{out}}=f_{\text{VE}}(x^{\prime}_{i})$ and $P^{\prime}_{\text{out}}=P(f_{\text{VE}}(x^{\prime}_{i}),x_{t})$ represent the corresponding outputs, where $x^{\prime}_{i}$ is initialized by adding random noise to $x_{i}$. We perform gradient-based adversarial attacks on both VE and VLP using the loss functions defined in Eqs. 2 and 3, where $I$ and $J$ denote the sequence length and feature dimension:
@@
The proposed loss function $\mathcal{L}_{\text{TCP}}$ combines these components, with $\beta\in[0,1]$ controlling the trade-off between VE and VLP losses, and $K$ denoting the number of surrogate VLPs used to enhance attack effectiveness. The overall loss function is:
$\mathcal{L}_{\text{TCP}}=\beta\mathcal{L}_{\text{VE}}+(1-\beta)\mathcal{L}_{\text{VLP}}$ (4)

### v2

@@ Abstract
The choice of a suitable visual language projector (VLP) is critical to the successful training of large visual language models (LVLMs). Mainstream VLPs can be broadly categorized into compressed and uncompressed projectors, and each offers distinct advantages in performance and computational efficiency. However, their security implications have not been thoroughly examined. Our comprehensive evaluation reveals significant differences in their security profiles: compressed projectors exhibit substantial vulnerabilities, allowing adversaries to successfully compromise LVLMs even with minimal knowledge of structure information. In stark contrast, uncompressed projectors demonstrate robust security properties and do not introduce additional vulnerabilities. These findings provide critical guidance for researchers in selecting optimal VLPs that enhance the security and reliability of visual language models. The code is available at https://github.com/btzyd/TCP.
The Security Threat of Compressed Projectors
in Large Vision-Language Models
Yudong Zhang1,2, Ruobing Xie2,✉, Xingwu Sun2,4, Jiansheng Chen3,✉, Zhanhui Kang2, Di Wang2, Yu Wang1,✉ 1Department of Electronic Engineering, Tsinghua University, 2Large Language Model Department, Tencent, 3School of Computer and Communication Engineering, University of Science and Technology Beijing, 4Faculty of Science and Technology, University of Macau zhangyd16@mails.tsinghua.edu.cn, xrbsnowing@163.com, sunxingwu01@gmail.com, jschen@ustb.edu.cn, kegokang@tencent.com, diwang@tencent.com, yu-wang@mail.tsinghua.edu.cn. (✉: Corresponding authors)
@@ 1 Introduction
Large vision-language models (LVLMs) have achieved remarkable success in various multimedia applications Dai et al. (2023); Liu et al. (2023a); Bai et al. (2025); Wang et al. (2024); Chen et al. (2024), particularly in tasks like visual question answering (VQA). A typical LVLM framework consists of three key components: visual encoder (VE), large language model (LLM), and vision-language projector (VLP). Current training methodologies primarily focus on optimizing the VLP to effectively integrate visual features extracted from a pre-trained VE into a pre-trained LLM. This approach substantially minimizes computational demands when compared to building an LVLM from scratch, thereby allowing researchers to leverage existing advancements. Moreover, it ensures greater stability and reliability in both the training process and practical deployment of LVLMs.
Recently, a growing body of research has employed CLIP Radford et al. (2021); Fang et al. (2023) as the VE and initiated training on various pre-trained LLMs. The choice of VLP plays a crucial role in determining the effectiveness of LVLM training. VLPs can generally be classified into two distinct categories: compressed projectors and uncompressed projectors. Compressed projectors, exemplified by Q-former Li et al. (2023); Dai et al. (2023); Zhu et al. (2024), achieve high computational efficiency by compressing a larger number of visual tokens into a smaller dimension using query tokens. On the other hand, uncompressed projectors, typified by MLP-based architectures Liu et al. (2023a); Shi et al. (2024), convert visual tokens to feature dimensions matching the LLM, with the number of output tokens proportional to the number of input tokens.
Previous studies have tended to focus only on the security of visual encoders or LLMs. On the visual encoder, Wang et al. (2023); Yin et al. (2023) present adversaries with an opportunity to compromise model security by targeting solely the visual encoder. As for LLMs, previous studies Carlini et al. (2023); Zou et al. (2023); Chao et al. (2023) have fully explored the safety of LLM, and the user has the option of choosing a better performing and safer LLM on his or her own. However, the security of visual language projectors has not been fully explored, and only a few studies have focused on attacks on VLPs. Although some studies Zhang et al. (2025) have attempted to attack the VLP (mainly Q-former), they have not thoroughly analyzed the security of the VLP structure.
@@
We assess the vulnerabilities of both compressed and uncompressed projectors by conducting adversarial attacks across diverse white-box and gray-box scenarios to evaluate the security implications for different VLP architectures. Our findings reveal three key insights: (1) Compressed projectors exacerbate security vulnerabilities beyond those affecting visual encoders. Specifically, an adversary can significantly degrade model performance by targeting compressed projectors, even with limited knowledge about the VLP. (2) In contrast, uncompressed projectors do not introduce additional security risks, as attacking them yields results comparable to attacking visual encoders directly. (3) This discrepancy in security between compressed and uncompressed projectors stems from their architectural differences and remains independent of the number of visual tokens. Notably, even when the visual tokens of an uncompressed projector are reduced (e.g., through pooling) to approach those of a compressed projector, the uncompressed projector still maintains its robust security characteristics.
Based on our analyses, we propose the following suggestions: (1) We strongly recommend employing uncompressed projectors rather than compressed projectors in high-security environments to mitigate potential risks from adversarial attacks. (2) In scenarios where computational efficiency is a priority, researchers can implement techniques like pooling operations to proportionally reduce the number of their output tokens. This approach offers enhanced security compared to using compressed projectors while maintaining acceptable performance levels, thereby striking an appropriate balance among effectiveness, efficiency, and security.
Our research provides three key contributions: (1) This study is the first to systematically investigate VLP security, presenting a comprehensive and systematic evaluation of various VLP architectures. (2) Our experimental results uncover significant security vulnerabilities in compressed projectors, revealing critical implications for models that employ such components. (3) We demonstrate that operating on visual tokens generated by uncompressed projectors offers enhanced security compared to compressed projectors when the number of visual tokens is largely comparable, thereby identifying promising approaches for VLP selection under efficiency constraints.
@@ 2 Related Work
Large Vision-Language Models. Training LVLMs from scratch is often prohibitively expensive. Thus, current popular frameworks typically begin with a pre-trained unimodal visual encoder and a large language model, focusing their training efforts on developing a VLP to connect the two modalities and effectively incorporate visual features into the LLM framework. The mainstream VLPs currently fall into two primary categories: compressed projectors and uncompressed projectors. Compressed projectors, such as Q-former Li et al. (2023), Resampler Alayrac et al. (2022), and D-Abstractor Cha et al. (2024), reduce the number of visual tokens to a fixed number of output tokens. In contrast, uncompressed projectors, exemplified by MLP Liu et al. (2023a), maintain a proportional relationship between the number of visual tokens and their corresponding output tokens.
Compressed and uncompressed projectors present distinct trade-offs in terms of computational efficiency, memory usage, and implementation complexity. On one hand, uncompressed projectors offer simplicity in design but incur significant computational costs. Specifically, Specifically, their computational requirements increase significantly as the token length scales linearly with the square of the resolution Chen et al. (2024). Furthermore, in video processing applications Ren et al. (2024), this scaling behavior results in a linear growth of required length relative to the number of video frames. On the other hand, compressed projectors achieve improved efficiency by reducing redundancy within the visual space. This is accomplished through the compression of token lengths from the visual encoder to a specified capacity, enabling strong performance while maintaining high computational efficiency. While prior research has extensively studied the performance and efficiency trade-offs between these two approaches in VLPs, there remains a critical gap: the safety implications of these methods have been largely neglected.
@@
This paper focuses on a practical gray-box scenario where the adversary is aware of both the VE weights and the structure of VLP, whether it be compressed or uncompressed. By crafting specialized loss functions tailored to the VLP architecture, the adversary can effectively execute targeted adversarial attacks against LVLMs in both white-box and gray-box settings. Our findings not only expose the vulnerabilities inherent in compressed projectors but also provide new insights into selecting between compressed and uncompressed projectors from a security perspective.
@@ 3.1 Preliminaries
Notations. We provide a brief overview of the definition of notations. An LVLM, denoted as $f$, generally comprises three key components: (1) the visual encoder $f_{\text{VE}}$, typically based on CLIP; (2) the vision-language projector $f_{\text{VLP}}$, which is usually implemented as a Q-former or an MLP; and (3) the large language model $f_{\text{LLM}}$. The LVLM $f$ accepts an image $x_{i}$ and an instruction $x_{t}$ as input, producing an output $y$. The process begins with the input image $x_{i}$ being processed by $f_{\text{VE}}$ to generate the visual feature representation $f_{\text{VE}}(x_{i})$. Next, depending on whether it is a compressed or uncompressed VLP: (1) for the compressed VLP, $f_{\text{VLP}}$ extracts relevant features from both the visual features $f_{\text{VE}}(x_{i})$ and the instruction $x_{t}$; while (2) for the uncompressed VLP, $f_{\text{VLP}}$ transforms the visual token dimension to align with the LLM input space. Both processes result in the projection output $f_{\text{VLP}}(f_{\text{VE}}(x_{i}),x_{t})$. These projected features are then fed into $f_{\text{LLM}}$, which generates the final output $y$ as Eq.˜1:
@@
Compressed projectors. The Q-former stands as the canonical representative of compressed projectors, functioning as a lightweight yet trainable querying transformer. It effectively extracts textually relevant features from the VE CLIP through a set of learnable queries. Notably, during BLIP-2’s two-stage pre-training process, both the VE and the LLM remain static, with only the Q-former and its queries being trainable components. This training paradigm has been successfully employed in subsequent studies, such as InstructBLIP Dai et al. (2023) and MiniGPT-4 Zhu et al. (2024), which also adopt similar approaches.
Uncompressed projectors. A typical approach involves using a simple multilayer perceptron (MLP) to project the visual token representation into the input space of an LLM. For instance, in LLaVA v1.5 Liu et al. (2023a), the visual encoder outputs a 1024-dimensional embedding, which is then mapped to 4096 dimensions via an MLP to align with the input requirements of the Vicuna LLM. Similarly, Eagle Shi et al. (2024) concatenates feature representations from multiple visual encoders to form a longer vector (typically ranging from 6000 to 8000 dimensions, depending on the number of visual encoders used), which is projected down to 4096 dimensions using an MLP.
@@ 3.2 An Empirical Analysis of Component Accessibility Within LVLM Attacks
Vision-language Projector (VLP). We assume the adversary possesses knowledge of the VLP’s architectural details, such as its implementation as a Q-former or MLP. Notably, all LVLMs built on Q-former, including BLIP-2, InstructBLIP, and MiniGPT-4, share identical VLP architectural designs. However, while the adversary may be aware of these structure specifics, obtaining the exact model parameters remains challenging due to the diversity in pre-trained LLMs and training datasets.
@@
The white-box setting. In the white-box setting, where adversaries have full access to model parameters (including both VEs and VLPs), we evaluate the safety of VLPs under worst-case scenarios. The adversary aims to modify a clean image to create an adversarial example capable of fooling the LVLM, leading to incorrect or unintended responses.
The gray-box setting. Building on the analysis of LVLMs’ components (VE, VLP, and LLM) presented earlier, we focus on a practical yet hitherto unexplored gray-box setting, in which the adversary has access to the weights of the VE CLIP and the architectural structure of the VLP but lacks access to either the weights of the VLP or any information about the LLM. In this context, we can obtain a surrogate VLP model through transfer learning from other LVLMs, or alternatively, train the surrogate VLP model from scratch.
@@ 3.3 Surrogate Models Used in Attacks
Based on our analysis of the accessibility levels of various LVLM components presented in Sec.˜3.2, we design and implement adversarial attacks against VLPs under three distinct scenarios with increasing levels of difficulty to comprehensively assess their security vulnerabilities. (1) The first scenario represents the simplest case, where the attack operates in a white-box setting, allowing full visibility into the model’s architecture and parameters. (2) The second scenario involves attacking a surrogate VLP from a similar LVLM, targeting another specific LVLM. (3) The most challenging scenario requires an attacker to develop a surrogate VLP from scratch, significantly increasing the complexity of generating effective adversarial examples.
@@
Surrogate models transferred from other LVLMs in gray-box setting. In our gray-box setting, while only the VE of the target model is available, we can still obtain surrogate VLPs by transferring knowledge from other similar LVLMs. For instance, if the target model is InstructBLIP Vicuna-13B, we can utilize the VLP from InstructBLIP $\text{FlanT5}_{\text{XL}}$ to attack it.
Surrogate models trained from scratch in gray-box setting. When similar LVLMs are unavailable for transfer, the structure insights into VLP provide us with a unique opportunity. Once the target model’s VLP structure (using compressed or uncompressed projectors) is identified, we can train surrogate Q-formers or MLPs from scratch. The detailed training procedure is outlined in Sec.˜4.1. Notably, during this process, no information about the specific LLM used by the target LVLM is required; only the VE weights and VLP structure are necessary for training the surrogate VLP.
Figure 1: The attack pipeline operates by first extracting surrogate VLPs $\{f_{\text{VLP}}^{1},\dots,f_{\text{VLP}}^{N}\}$, which are then utilized to generate adversarial examples $x^{\prime}_{i}$ through the loss functions $\mathcal{L}_{\text{VE}}$, $\mathcal{L}_{\text{VLP}}$ and $\mathcal{L}_{\text{TCP}}(\beta,K)$ (TCP stands for “Threat of Compressed Projectors”). A key aspect of this investigation is determining whether incorporating attacks on VLPs increases security vulnerabilities compared to solely attacking VEs, thereby assessing the robustness of the VLP structure.
@@ 3.4 Loss Function for More Effective Attacks
To evaluate the security of the VLP structure, we employ adversarial attack methods to analyze the robustness of LVLM. For a model with loss function $\mathcal{L}$, Fast Gradient Sign Method (FGSM) Goodfellow et al. (2015) performs an adversarial attack through gradient ascent as follows: $x^{\prime}_{i}=x^{\prime}_{i}+\nabla_{x^{\prime}_{i}}\mathcal{L}$. In contrast, Projected Gradient Descent (PGD) Madry et al. (2018) conducts multiple iterations of gradient ascent and projects $x^{\prime}_{i}$ onto the constrained perturbation space after each step. Typically, PGD restricts the $l_{\infty}$ norm distance with $\|x_{i}-x^{\prime}_{i}\|_{\infty}\leq\epsilon$. Additionally, a variant of the Carlini & Wagner (C&W) attack Carlini and Wagner (2017), referred to as CW-$l_{2}$, aims to maximize the loss function $\mathcal{L}$ while minimizing the $l_{2}$ norm distance $x_{i}-x^{\prime}_{i}||_{2}^{2}$.
To compare the security implications introduced by the VLP structure, we evaluate the attack results on VE and VLP. For a clean image input, let $V_{\text{out}}=f_{\text{VE}}(x_{i})$ and $P_{\text{out}}=P(f_{\text{VE}}(x_{i}),x_{t})$ denote the outputs of VE and VLP, respectively. For an adversarial image, let $V^{\prime}_{\text{out}}=f_{\text{VE}}(x^{\prime}_{i})$ and $P^{\prime}_{\text{out}}=P(f_{\text{VE}}(x^{\prime}_{i}),x_{t})$ represent the corresponding outputs, where $x^{\prime}_{i}$ is initialized by adding random noise to $x_{i}$. We perform gradient-based adversarial attacks on both VE and VLP using the loss functions defined in Eqs.˜2 and 3, where $I$ and $J$ denote the sequence length and feature dimension:
@@
The proposed loss function $\mathcal{L}_{\text{TCP}}$ (TCP stands for “Threat of Compressed Projectors”) combines these components, with $\beta\in[0,1]$ controlling the trade-off between VE and VLP losses, and $K$ denoting the number of surrogate VLPs used to enhance attack effectiveness. The overall loss function is:
$\mathcal{L}_{\text{TCP}}=\beta\mathcal{L}_{\text{VE}}+(1-\beta)\times\frac{1}{K}\Sigma_{j=1}^{K}\mathcal{L}_{\text{VLP}}^{j}$ (4)